OpenStudy (anonymous):

verify the following identities. HELP PLEASE ! 2. sec^2xcotx-cotx=tanx 3.sinxcosxtanx=1-cosx 4.tanx+cotx=secxcscx

OpenStudy (anonymous):

2. \[\sec^{2}x = 1+ \tan^{2}x\] \[cotx = \frac{1}{tanx}\] hence, \[[(1+\tan^{2}x)(\frac{1}{tanx})] - \frac{1}{tanx}\]

OpenStudy (anonymous):

you get \[tanx + \frac{1}{tanx} -\frac{1}{tanx} = tanx\]

OpenStudy (anonymous):

for 3. \[tanx = \frac{sinx}{cosx}\] Sub, factor and solve.. for 4. \[secx= \frac{1}{cosx}, cscx= \frac{1}{sinx}\]

OpenStudy (anonymous):

If you can get that on the left side for 4 your good.

OpenStudy (anonymous):

how did you get that on 4?

OpenStudy (anonymous):

get what?